### v1

# Generative Modeling through Spectral Analysis of Koopman Operator

Abstract
We propose Koopman Spectral Wasserstein Gradient Descent (KSWGD), a generative modeling framework that combines operator-theoretic spectral analysis with optimal transport. The novel insight is that the spectral structure required for accelerated Wasserstein gradient descent can be directly estimated from trajectory data via Koopman operator approximation which can eliminate the need for explicit knowledge of the target potential or neural network training. We provide rigorous convergence analysis and establish connection to Feynman-Kac theory that clarifies the method’s probabilistic foundation. Experiments across diverse settings, including compact manifold sampling, metastable multi-well systems, image generation, and high dimensional stochastic partial differential equation, demonstrate that KSWGD consistently achieves faster convergence than other existing methods while maintaining high sample quality.
Keywords Koopman operator $\cdot$ Dynamic mode decomposition $\cdot$ Spectral analysis $\cdot$ Langevin dynamics $\cdot$ Wasserstein gradient flow $\cdot$ Feynman-Kac formula $\cdot$ Generative modeling
1 Introduction
Generative modeling plays a central role in modern data science, with applications ranging from computer vision and natural language processing to scientific computing and uncertainty quantification [arjovsky2017wasserstein, baikonode, cheng2025safe, dinh2017realnvp, geng2025mean, goodfellow2014generative, ho2020denoising, nichol2021improved, radford2018improving, sohl2015deep, song2019generative, song2020score, van2016pixel, yu2025unlock]. Given samples from an unknown target distribution, the goal is to generate additional samples without explicit access to its density. Despite their empirical success, many popular generative models, such as variational autoencoders [kingma2013auto, rezende2014stochastic], generative adversarial networks [goodfellow2014generative], and diffusion or score-based models [ho2020denoising, song2019generative, song2020score, yang2023diffusion], rely on extensive neural network training, careful hyperparameter tuning, and substantial computational resources. Moreover, Langevin-type sampling and score-based approaches require access to gradients of the log-density or learned score functions, which can be costly or unstable to estimate in high-dimensional or complex settings.
An alternative perspective on sampling and generative modeling is provided by Wasserstein gradient flows [ambrosio2008gradient, arbel2019maximum, fan2021variational, jordan1998variational, liu2017stein, liu2016stein, mokrov2021large], which offer a unifying variational framework for many stochastic and deterministic particle methods. Algorithms such as Stein variational gradient descent [liu2017stein, liu2016stein], affine-invariant Langevin dynamics [garbuno2020affine], and Laplacian-adjusted Wasserstein gradient descent (LAWGD) [chewi2020svgd] can be interpreted as discrete approximation of gradient flows on the space of probability measures. In particular, LAWGD achieves strong theoretical guarantees by preconditioning the gradient flow with the inverse Langevin generator $\mathcal{L}^{-1}$. However, the inverse generator is typically inaccessible due to the unknown target density.
Recent work has therefore focused on data-driven spectral approximations of the Langevin generator. A notable example is the diffusion map particle system (DMPS) [li2025diffusion], which combines diffusion maps with LAWGD to construct a kernel-based approximation of the generator from samples. Diffusion maps [coifman2006diffusion, coifman2005geometric, coifman2008diffusion, nadler2006diffusion] construct a kernel integral operator based on pairwise distances that provides a consistent approximation of the overdamped Langevin generator $\mathcal{L}$ under appropriate density normalization. This nonparametric construction is training-free and is particularly effective for distributions supported on low-dimensional manifolds. However, diffusion maps only use pairwise distances between i.i.d. samples and require careful tuning of the kernel bandwidth to approximate the generator accurately. Temporal ordering and trajectory information are not explicitly used in the operator construction, which limits the direct applicability of such methods to time-series data or prediction tasks arising from dynamical systems.
In this work, we choose an operator-theoretic perspective grounded in the Kolmogorov backward equation [engel1999one, oksendal2003stochastic], which governs the evolution of conditional expectations under stochastic dynamics. For overdamped Langevin systems, the infinitesimal generator of this evolution coincides with the Langevin generator, i.e., $\mathcal{A}=$-$\mathcal{L}$, up to a negative sign, and serves as the Koopman generator of the underlying Markov semigroup [engel1999one]. This connection allows the generator to be approximated directly from trajectory data using spectral methods such as extended dynamic mode decomposition (EDMD) [Williams_2015], kernel EDMD [kevrekidis2016kernel], and neural network-based Koopman approaches [lusch2018deep, li2017extended, xu2025reskoopnet]. While diffusion maps approximate the generator based on static pairwise geometry, Koopman-based methods can also explicitly incorporate temporal evolution through time-ordered snapshot pairs. Using Galerkin projection onto a finite-dimensional dictionary space, the spectral approximation enables the construction of the inverse operator $\mathcal{L}^{-1}$ required by LAWGD. Unlike general applications of Koopman operator theory to chaotic or complex systems, where the spectrum is non-real and eigenfunctions are non-orthogonal, the generative modeling task relies on Langevin dynamics, which satisfies the detailed balance condition. This physical property induces a crucial mathematical structure that the associated Koopman generator is self-adjoint (Hermitian) with respect to the invariant measure $\pi$. The self-adjointness guarantees that the spectrum is purely real and the eigenfunctions form an orthogonal basis of $L^{2}_{\pi}$, which is an important mathematical property and valuable to us.
Building on this insight, we propose Koopman Spectral Wasserstein Gradient Descent (KSWGD), a training-free generative modeling framework that integrates Wasserstein gradient flows with modern Koopman spectral analysis [brunton2022data, brunton2021modern, mezic2005spectral]. Replacing the inaccessible inverse Langevin generator with its Koopman-based spectral approximation leads to deterministic particle dynamics with a constant dissipation rate, preventing the vanishing-gradient behavior observed in kernel-based particle methods [duncan2023geometry, fujisawa2025on, liu2023towards, wainwright2019high]. We establish linear convergence in discrete time, together with explicit error bounds determined by spectral truncation and operator approximation accuracy. The Koopman formulation further admits a natural linear Feynman–Kac interpretation [oksendal2003stochastic], which provides a rigorous probabilistic foundation for unconditional sampling.
The main contributions of this paper are:
  1. 1.
    We proposed a training-free generative modeling framework based on Koopman spectral approximations of the inverse Langevin generator.
  2. 2.
    Convergence guarantees with explicit error bounds arising from spectral truncation and numerical approximation.
  3. 3.
    A (backward) operator-theoretic perspective that unifies Wasserstein gradient flows, Koopman theory, and generative particle systems.
  4. 4.
    Empirical validation of data in the form of both time series and static time on compact manifold, metastable systems, image data, and high-dimensional stochastic partial differential equation.
The remainder of the paper is organized as follows. Section 2 reviews the operator-theoretic background on Wasserstein gradient flows and Koopman semigroups. Section 3 introduces the proposed Koopman Spectral Wasserstein Gradient Descent (KSWGD) framework and its spectral construction. Section 4 establishes convergence guarantees and error bounds based on numerical spectral truncation. Section 5 provides a Feynman–Kac interpretation of KSWGD, clarifying its connection to unconditional sampling and potential extension to conditional sampling problems. Numerical experiments are presented in Section 6, followed by conclusion and future directions in Section 7.
2 Preliminary
2.1 Wasserstein Gradient Flow
Let $\pi(x)\propto e^{-V(x)}$ be a target distribution on $\mathbb{R}^{d}$. The Jordan–Kinderlehrer–Otto (JKO) framework [jordan1998variational] interprets evolution equations for probability measures as Wasserstein gradient flows on $(\mathcal{P}_{2}(\mathbb{R}^{d}),W_{2})$ where $W_{2}$ denotes the $2$-Wasserstein distance [santambrogio2015optimal].
For a functional $\mathcal{F}:\mathcal{P}_{2}(\mathbb{R}^{d})\to\mathbb{R}$, the Wasserstein gradient flow is defined through the continuity equation
$\partial_{t}\mu_{t}=\operatorname{div}(\mu_{t}\,\nabla_{W_{2}}\mathcal{F}(\mu_{t})),$
where $\nabla_{W_{2}}\mathcal{F}(\mu):\mathbb{R}^{d}\to\mathbb{R}^{d}$ denotes the Wasserstein gradient [santambrogio2015optimal]. This PDE describes the evolution of the distribution $\mu_{t}$ whose law is governed by particles moving along the velocity field
$v_{t}(x)=-\nabla_{W_{2}}\mathcal{F}(\mu_{t})(x).$
Classical examples include the Kullback–Leibler (KL) divergence $\mathrm{KL}(\mu\|\pi)\coloneqq\int\log(\mathrm{d}\mu/\mathrm{d}\pi)\,d\mu$ with Wasserstein gradient $\nabla_{W_{2}}\mathrm{KL}(\mu)=\nabla\log(\mathrm{d}\mu/\mathrm{d}\pi)$, and the chi-squared divergence $\chi^{2}(\mu\|\pi)\coloneqq\mathbb{E}_{\pi}[(\mathrm{d}\mu/\mathrm{d}\pi-1)^{2}]$ with $\nabla_{W_{2}}\chi^{2}(\mu)=2\nabla(\mathrm{d}\mu/\mathrm{d}\pi)$ [ambrosio2008gradient, santambrogio2015optimal, Tsybakov2009]. The chi-squared gradient flow enables sharper convergence analysis for deterministic particle methods [chewi2020svgd].
LAWGD. The chi-squared gradient flow $\partial_{t}\mu_{t}=2\operatorname{div}(\mu_{t}\nabla(\mathrm{d}\mu_{t}/\mathrm{d}\pi))$ admits strong theoretical guarantees but is computationally intractable, in other words, evaluating the velocity field $2\nabla(\mathrm{d}\mu_{t}/\mathrm{d}\pi)$ requires estimating the density ratio $\mathrm{d}\mu_{t}/\mathrm{d}\pi$ at each iteration, which suffers from the curse of dimensionality. To circumvent this difficulty, Laplacian-Adjusted Wasserstein Gradient Descent (LAWGD) [chewi2020svgd] applies a spectral preconditioning via the integral operator $\mathcal{K}_{\pi}$:
$\partial_{t}\mu_{t}=\operatorname{div}\!\left(\mu_{t}\,\nabla\mathcal{K}_{\pi}(\mathrm{d}\mu_{t}/\mathrm{d}\pi)\right),$ (2.1)
where $\mathcal{K}_{\pi}f(x)\coloneqq\int K(x,y)f(y)\,\mathrm{d}\pi(y)$ is the integral operator associated with a carefully chosen kernel $K$. The most important design principle here is to select $K$ such that $\mathcal{K}_{\pi}=\mathcal{L}^{-1}$ where $\mathcal{L}\coloneqq-\Delta+\langle\nabla V,\nabla\cdot\rangle$ is the Langevin generator associated with $\pi$ and the inverse is understood as the inverse of $\mathcal{L}$ restricted to the orthogonal complement of constant functions in $L^{2}_{\pi}$. Notice that, here $\mathcal{L}$ is a positive self-adjoint operator on $L^{2}_{\pi}$. For standard distribution with exponential or faster tail decay, $\mathcal{L}$ has a discrete spectrum $0=\lambda_{0}<\lambda_{1}\leq\lambda_{2}\leq\cdots$ and corresponding orthonormal eigenfunctions $\{\phi_{i}\}_{i\geq 0}$ in $L^{2}_{\pi}$ with $\phi_{0}\equiv 1$ (constant function). The LAWGD kernel is then constructed via spectral representation:
$K_{\mathcal{L}^{-1}}(x,y)\coloneqq\sum_{i=1}^{\infty}\frac{\phi_{i}(x)\phi_{i}(y)}{\lambda_{i}},$ (2.2)
where the summation excludes the zero eigenvalue to ensure well-definedness. This choice yields a remarkable dissipation identity. To see this, note that by integration by parts [chewi2020svgd, Eq. (10)]:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\mu_{t}\|\pi)=-\left\langle\nabla\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\pi},\nabla\mathcal{K}_{\pi}\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\pi}\right\rangle_{\pi}=-\left\langle\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\pi}-1,\mathcal{L}\mathcal{K}_{\pi}\left(\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\pi}-1\right)\right\rangle_{\pi}=-\chi^{2}(\mu_{t}\|\pi),$ (2.3)
where the second equality uses the Dirichlet form $\langle\nabla f,\nabla g\rangle_{\pi}=\langle f,\mathcal{L}g\rangle_{\pi}$. This constant dissipation rate implies scale-free exponential convergence $\mathrm{KL}(\mu_{t}\|\pi)\leq\mathrm{KL}(\mu_{0}\|\pi)e^{-t}$ independent of the Poincaré constant [chewi2020svgd, Theorem 4]. Particle implementation. The continuity equation (2.1) translates to particle dynamics via the velocity field $v_{t}(x)=-\nabla\mathcal{K}_{\pi}(\mathrm{d}\mu_{t}/\mathrm{d}\pi)(x)$. Using the identity
$\nabla\mathcal{K}_{\pi}f(x)=\int\nabla_{1}K(x,y)f(y)\,\mathrm{d}\pi(y)=\int\nabla_{1}K_{\mathcal{L}^{-1}}(x,y)\,\mathrm{d}\mu_{t}(y),$
where $\nabla_{1}$ denotes gradient w.r.t. the first argument, particles $\{x_{t}^{(i)}\}_{i=1}^{M}$ evolve according to
$\dot{x}_{t}^{(i)}=-\frac{1}{M}\sum_{j=1}^{M}\nabla_{1}K_{\mathcal{L}^{-1}}(x_{t}^{(i)},x_{t}^{(j)}),\;1\leq i\leq M,$ (2.4)
where $\mu_{t}\approx\frac{1}{M}\sum_{j=1}^{M}\delta_{x_{t}^{(j)}}$ is the empirical measure.
Towards data-driven implementation. The implementation of LAWGD relies on the spectral decomposition of the Langevin generator $\mathcal{L}$, which requires explicit access to the potential function $V$. In many applications, however, $V$ is unknown or computationally inaccessible, i.e., the target distribution $\pi$ may be specified only through samples, or the dynamics may arise from complex systems where no closed-form potential exists. This gap between theory and practice motivates the development of data-driven methods for spectral approximation.
2.2 Koopman Operator
Consider a dynamical system governed by the SDE
$\mathrm{d}X_{t}=b(X_{t})\mathrm{d}t+\sigma(X_{t})\mathrm{d}W_{t},$ (2.5)
where $b:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the drift, $\sigma:\mathbb{R}^{d}\to\mathbb{R}^{d\times m}$ is the diffusion coefficient satisfying appropriate regularity condition [engel1999one], and $W_{t}$ is an $m$-dimensional Brownian motion.
Koopman operator theory [Koopman1931, koopman1932dynamical, mezic2005spectral] provides a linear framework for analyzing such systems by lifting the nonlinear dynamics to a linear operator acting on observable functions. For the continuous time stochastic system above, the Koopman operator $\mathcal{T}^{t}:L^{2}(\mathcal{X})\to L^{2}(\mathcal{X})$ acts on observable functions $g:\mathcal{X}\to\mathbb{R}$ via
$(\mathcal{T}^{t}g)(x)\coloneqq\mathbb{E}[g(X_{t})\mid X_{0}=x],$ (2.6)
where the conditional expectation is taken over all realizations of the Brownian motion given the initial state $x$. This operator is linear despite the potential nonlinearity of $b$ and $\sigma$, enabling the application of spectral methods to analyze the stochastic dynamics. The infinitesimal generator $\mathcal{A}$ of the Koopman operator is defined as
$\mathcal{A}f\coloneqq\lim_{t\to 0}\frac{\mathcal{T}^{t}f-f}{t},$
on the domain
$\mathcal{D}(\mathcal{A})=\left\{f\in L^{2}(\mathcal{X}):\lim_{t\to 0}\frac{\mathcal{T}^{t}f-f}{t}\text{ exists in }L^{2}(\mathcal{X})\right\}.$
By Itô’s formula [engel1999one], this generator has the following explicit form
$\mathcal{A}f(x)=\langle b(x),\nabla f(x)\rangle+\frac{1}{2}\text{Tr}\left(\sigma(x)\sigma(x)^{\top}\nabla^{2}f(x)\right),$ (2.7)
for all $x\in\mathcal{X}$ and $f\in C^{2}_{b}(\mathcal{X})\cap L^{2}(\mathcal{X})$.
In practice, the eigenpairs of the Koopman operator can be directly estimated from data. Several data-driven methods have been developed for Koopman operator approximation in deterministic or stochastic settings such as EDMD [Williams_2015] and etc. [arbabi2017ergodic, baikonode, cheng2025information, colbrook2024rigorous, gao2025conformalonlinelearningdeep, ishikawa2024koopman, kevrekidis2016kernel, kostic2022learning, mauroy2024analytic, oprea2025distributional, xu2025reinforced, 10.1063/5.0283640, xu2025reskoopnet, pmlr-v267-zhang25be].
3 Data-Driven Spectral Construction
3.1 Koopman Interpretation of Wasserstein Gradient Descent
While LAWGD [chewi2020svgd] provides a theoretically appealing sampling dynamics with scale-free exponential convergence, its formulation assumes access to the exact inverse of the Langevin generator. More precisely, for the overdamped Langevin dynamics, i.e., letting $b=-\nabla V$ and $\sigma=\sqrt{2}I_{d}$ for (2.5), the Langevin generator [pavliotis2016stochastic] is $\mathcal{L}=-\Delta+\langle\nabla V,\nabla\cdot\rangle$ [pavliotis2016stochastic]. In this idealized setting, the algorithm requires the full spectral decomposition of $\mathcal{L}$, and thus it is not directly implementable when $V$ or $\mathcal{L}$ are unknown.
A key observation is that the Langevin generator $\mathcal{L}$ coincides (up to a sign) with the Koopman generator of the Markov semigroup associated with $(X_{t})_{t\geq 0}$. In other words, if $(\mathcal{T}^{t})_{t\geq 0}$ denotes the Koopman semigroup, then for any smooth test function $f$, its infinitesimal generator $\mathcal{A}$ satisfies
$\mathcal{A}f(x)=-\langle\nabla V(x),\nabla f(x)\rangle+\Delta f(x)=-\mathcal{L}f(x),$
by (2.7). Hence $\mathcal{A}=$-$\mathcal{L}$ on the common domain of definition. This link makes Koopman spectral methods a natural tool for approximating the eigenpairs of $\mathcal{L}$ from data. Since LAWGD requires access to $\mathcal{L}^{-1}$, the Koopman spectral decomposition provides a natural finite-dimensional surrogate through the first $r$ eigenpairs $\{(\lambda_{i},\phi_{i})\}_{i=1}^{r}$. We define the truncated inverse operator
$\mathcal{K}_{r}\coloneqq\sum_{i=1}^{r}\frac{1}{\lambda_{i}}\,\langle\cdot,\phi_{i}\rangle_{\pi}\,\phi_{i},$ (3.1)
which recovers the exact inverse, i.e., $\mathcal{L}^{-1}$, as $r\to\infty$ as shown in [10.1063/5.0283640] and plays the role of the LAWGD kernel in the Koopman-based construction.
3.2 Truncated Koopman Approximation and Particle Dynamics
In practice, we approximate the leading eigenpairs $\{(\lambda_{i},\phi_{i})\}_{i=1}^{r}$ of $\mathcal{L}$ through data-driven estimators of the Koopman generator $\mathcal{A}$, e.g., EDMD [Williams_2015] or kernel-EDMD [kevrekidis2016kernel]. Given training samples $\{z^{(j)}\}_{j=1}^{N}\sim\pi$, these methods construct a data-driven approximation $\widehat{\mathcal{A}}_{N}$ on a $N$-dimensional subspace whose leading $r$ eigenpairs $\{(\widehat{\lambda}_{i},\widehat{\phi}_{i})\}_{i=1}^{r}$ give the $r$-dimensional approximated inverse operator
$\widehat{\mathcal{K}}_{r}=\sum_{i=1}^{r}\frac{1}{\widehat{\lambda}_{i}}\langle\cdot,\widehat{\phi}_{i}\rangle_{\pi}\widehat{\phi}_{i}.$ (3.2)
This construction defines a Koopman-based approximation of the integral operator $\mathcal{K}_{\pi}$ constructed by the LAWGD kernel as in (2.2), preserving the functional form of the LAWGD dynamics while introducing spectral truncation from the finite rank $r$.
Next, the particle dynamics require evaluating the gradient of the truncated kernel
$K_{\widehat{\mathcal{K}}_{r}}(x,y)=\sum_{i=1}^{r}\frac{\widehat{\phi}_{i}(x)\widehat{\phi}_{i}(y)}{\widehat{\lambda}_{i}}.$
For each particle pair $(x^{(i)}_{t},x^{(j)}_{t})$, we compute
$\nabla_{1}K_{\widehat{\mathcal{K}}_{r}}(x^{(i)}_{t},x^{(j)}_{t})=\sum_{k=1}^{r}\frac{\nabla\widehat{\phi}_{k}(x^{(i)}_{t})\cdot\widehat{\phi}_{k}(x^{(j)}_{t})}{\widehat{\lambda}_{k}},$ (3.3)
where $\nabla_{1}$ denotes the gradient with respect to the first argument and $\nabla\widehat{\phi}_{k}$ can be computed analytically for smooth basis functions (e.g., Gaussian kernels) or approximated via finite differences. Following (2.4), the Koopman-adjusted particle update becomes
$x_{t+1}^{(i)}=x_{t}^{(i)}-\frac{h}{M}\sum_{j=1}^{M}\nabla_{1}K_{\widehat{\mathcal{K}}_{r}}(x_{t}^{(i)},x_{t}^{(j)}),\quad 1\leq i\leq M.$ (3.4)
Input: Training samples $\{z^{(j)}\}_{j=1}^{N}\sim\pi$, initial particles $\{x_{0}^{(i)}\}_{i=1}^{M}$, step size $h$, truncation rank $r$, dictionary size $n$, max iterations $T$.
1
21exConstruct dictionary $\{\psi_{k}\}_{k=1}^{n}$ and estimate Koopman operator using trajectory pairs $\{(z^{(j)},z^{(j+1)})\}_{j=1}^{N-1}$ ;
3 Compute leading $r$ eigenpairs $\{(\widehat{\lambda}_{i},\widehat{\phi}_{i})\}_{i=1}^{r}$ of $\widehat{\mathcal{A}}_{N}$ ;
4
51exfor $t=0,1,\ldots,T-1$ do
6 for $i=1,\ldots,M$ do
7    $\mathbf{v}_{i}\leftarrow\mathbf{0}\in\mathbb{R}^{d}$ ;
8    for $j=1,\ldots,M$ do
9       $\mathbf{v}_{i}\leftarrow\mathbf{v}_{i}+\sum_{k=1}^{r}\frac{\nabla\widehat{\phi}_{k}(x_{t}^{(i)})\cdot\widehat{\phi}_{k}(x_{t}^{(j)})}{\widehat{\lambda}_{k}}$ by Eq. (3.3)
10    $x_{t+1}^{(i)}\leftarrow x_{t}^{(i)}-\frac{h}{M}\mathbf{v}_{i}$ by Eq. (3.4)
11
Output: Generated particles $\{x_{T}^{(i)}\}_{i=1}^{M}$.
Algorithm 1 KSWGD for Time Series Data
Input: Training samples $\{z^{(j)}\}_{j=1}^{N}\sim\pi$, initial particles $\{x_{0}^{(i)}\}_{i=1}^{M}$, step size $h$, time step $\Delta t$, KDE bandwidth $\sigma$, truncation rank $r$, dictionary size $n$, max iterations $T$.
1
21exEstimate score: $\nabla\log\widehat{\pi}(x)$ using KDE ;
3 for $j=1,\ldots,N$ do
4   Sample $\xi^{(j)}\sim\mathcal{N}(0,I_{d})$ ;
5 $\hat{z}_{\Delta t}^{(j)}\leftarrow z^{(j)}-\Delta t\cdot\nabla\log\widehat{\pi}(z^{(j)})+\sqrt{2\Delta t}\,\xi^{(j)}$ ;
6
7
81exConstruct dictionary $\{\psi_{k}\}_{k=1}^{n}$ and estimate Koopman operator using pairs $\{(z^{(j)},\hat{z}_{\Delta t}^{(j)})\}_{j=1}^{N}$ ;
9 Compute leading $r$ eigenpairs $\{(\widehat{\lambda}_{i},\widehat{\phi}_{i})\}_{i=1}^{r}$ of $\widehat{\mathcal{A}}_{N}$ ;
10
111exfor $t=0,1,\ldots,T-1$ do
12 for $i=1,\ldots,M$ do
13    $\mathbf{v}_{i}\leftarrow\mathbf{0}\in\mathbb{R}^{d}$ ;
14    for $j=1,\ldots,M$ do
15       $\mathbf{v}_{i}\leftarrow\mathbf{v}_{i}+\sum_{k=1}^{r}\frac{\nabla\widehat{\phi}_{k}(x_{t}^{(i)})\cdot\widehat{\phi}_{k}(x_{t}^{(j)})}{\widehat{\lambda}_{k}}$
16    $x_{t+1}^{(i)}\leftarrow x_{t}^{(i)}-\frac{h}{M}\mathbf{v}_{i}$
17
Output: Generated particles $\{x_{T}^{(i)}\}_{i=1}^{M}$
Algorithm 2 KSWGD for Static Time Data
Algorithm 1 and  2 summarized the complete procedure for time series type of data and static time data (e.g., image dataset). The computational cost is dominated by the offline eigendecomposition ($O(n^{3})$ for basis size $n$) and the online kernel gradient evaluation ($O(M^{2}rd)$ per iteration for $M$ particles in dimension $d$ with rank $r$). Unlike score-based methods that require neural network training, KAWGD’s sampling phase is deterministic and training-free once the Koopman basis is determined. Notice that the gradient of eigenfunctions in (3.3) would be difficult to compute. In this work, we compute it by Diffusion map with Gaussian RBF kernel. See [li2025diffusion] for more details.
4 Convergence and Error Bound Analysis
4.1 Data-Driven Error Bound Analysis
In this section, we analyze the convergence properties of the Koopman spectral Wasserstein gradient descent dynamics introduced in Section 3. Throughout the section, we denote
$\rho_{t}\coloneqq\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\pi},\qquad f_{t}\coloneqq\rho_{t}-1\in L_{0}^{2}(\pi),$
the fluctuation of the density with respect to $\pi$. Let $\{(\phi_{i},\lambda_{i})\}_{i\geq 1}$ be the eigenpairs of the Langevin generator $\mathcal{L}$, and recall the truncated inverse operator in (3.1), we have
$\mathcal{L}\mathcal{K}_{r}=\Pi_{r},$ (4.1)
where $\Pi_{r}:L^{2}_{\pi}\to\text{span}\{\phi_{1},\ldots,\phi_{r}\}$ denotes the orthogonal projector, and $\{\phi_{i}\}_{i\geq 1}$ are the exact $L^{2}_{\pi}$-orthonormal eigenfunctions of $\mathcal{L}$. To quantify the contribution of unretained high-order modes, we define the spectral tail error
$\eta_{r}(f):=\|(I-\Pi_{r})f\|_{L^{2}_{\pi}}=\left(\sum_{i>r}\langle f,\phi_{i}\rangle_{\pi}^{2}\right)^{1/2}.$
Assumption 4.1 (Regularity).
For regularity conditions, we assume that $\rho_{t}>0\textit{ a.e.}$ and $\partial_{t}(\rho_{t}\log\rho_{t})\in L^{1}_{\pi}$. In addition, we assume that $\nabla f_{t}\in L^{2}(\pi)$ and $\mathcal{K}_{r}f_{t}\in\mathcal{D}(\mathcal{L})$.
Assumption 4.2 (Uniform spectral tail bound).
There exists a constant $\eta_{r}>0$ such that
$\sup_{t\geq 0}\|(I-\Pi_{r})f_{t}\|_{L^{2}_{\pi}}\;\leq\;\eta_{r}.$
Proposition 4.3 (Convergence with spectral truncation).
Assume that Assumptions 4.1 and 4.2 holds. Let $(\mu_{t})_{t\geq 0}$ be the solution to the truncated LAWGD dynamics with exact eigenpairs and initial distribution $\mu_{0}$:
$\partial_{t}\mu_{t}=\operatorname{div}\big(\mu_{t}\nabla\mathcal{K}_{r}(\mathrm{d}\mu_{t}/\mathrm{d}\pi)\big),$
where $\mathcal{K}_{r}=\sum_{i=1}^{r}\frac{1}{\lambda_{i}}\langle\cdot,\phi_{i}\rangle_{\pi}\phi_{i}$ defined in (3.1) satisfies (4.1). Then
$\mathrm{KL}(\mu_{t}\|\pi)\leq\mathrm{KL}(\mu_{0}\|\pi)e^{-t}+\eta_{r}^{2}(1-e^{-t}).$ (4.2)
Proof.
Let $\rho_{t}=\frac{\mathrm{d}\mu_{t}}{\mathrm{d}\pi}$ and $f_{t}=\rho_{t}-1$. Recall the KL divergence $\mathrm{KL}(\mu_{t}\|\pi)=\int\rho_{t}\log\rho_{t}\,\mathrm{d}\pi$. By differentiation under the invariant measure $\pi$, see [chewi2023log, chewi2020svgd] for more details), we obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\mu_{t}\|\pi)$ $=\int\partial_{t}(\rho_{t}\log\rho_{t})\,\mathrm{d}\pi$
$=\int\left[(\partial_{t}\rho_{t})\log\rho_{t}+\rho_{t}\partial_{t}(\log\rho_{t})\right]\,\mathrm{d}\pi$
$=\int\mathrm{div}\!\bigl(\rho_{t}\,\nabla\mathcal{K}_{r}f_{t}\bigr)\,\log\rho_{t}\,\mathrm{d}\pi+\int\partial_{t}\rho_{t}\,\mathrm{d}\pi$
$=-\int\big\langle\nabla\log\rho_{t},\,\nabla\mathcal{K}_{r}f_{t}\big\rangle\,\rho_{t}\,\mathrm{d}\pi$
$=-\big\langle\nabla\log\rho_{t},\,\nabla\mathcal{K}_{r}f_{t}\big\rangle_{\mu_{t}},$
where $\int\partial_{t}\rho_{t}\,\mathrm{d}\pi=\frac{\mathrm{d}}{\mathrm{d}t}\int\rho_{t}\,\mathrm{d}\pi=0$. Then, using $\rho_{t}=1+f_{t}$ we have the identity
$\big\langle\nabla\log\rho_{t},\,\nabla g\big\rangle_{\mu_{t}}=\int\Big\langle\frac{\nabla\rho_{t}}{\rho_{t}},\,\nabla g\Big\rangle\rho_{t}\,\mathrm{d}\pi=\int\langle\nabla\rho_{t},\,\nabla g\rangle\,\mathrm{d}\pi=\int\langle\nabla f_{t},\,\nabla g\rangle\,\mathrm{d}\pi=\big\langle\nabla f_{t},\,\nabla g\big\rangle_{\pi}.$
Applying this with $g=\mathcal{K}_{r}f_{t}$ gives
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\mu_{t}\|\pi)=-\big\langle\nabla f_{t},\,\nabla\mathcal{K}_{r}f_{t}\big\rangle_{\pi}=-\langle f_{t},\mathcal{L}\mathcal{K}_{r}f_{t}\rangle_{\pi}=-\langle f_{t},\Pi_{r}f_{t}\rangle_{\pi},$ (4.3)
where we used the Dirichlet form identity $\langle\nabla f,\nabla g\rangle_{\pi}=\langle f,\mathcal{L}g\rangle_{\pi}$ and the operator relation (4.1).
Expanding the right-hand side of (4.3), we have
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\mu_{t}\|\pi)=-\langle f_{t},\Pi_{r}f_{t}\rangle_{\pi}=-\|\Pi_{r}f_{t}\|_{L^{2}_{\pi}}^{2},$ (4.4)
since $\Pi_{r}$ is the orthogonal projection, which is self-adjoint and idempotent and thus implies $\langle f_{t},\Pi_{r}f_{t}\rangle_{\pi}=\|\Pi_{r}f_{t}\|_{L^{2}_{\pi}}^{2}$.
If, in addition, $\|(I-\Pi_{r})f_{t}\|_{L^{2}_{\pi}}\leq\eta_{r}$, then
$\|f_{t}\|_{L^{2}_{\pi}}^{2}=\|\Pi_{r}f_{t}\|_{L^{2}_{\pi}}^{2}\;+\;\|(I-\Pi_{r})f_{t}\|_{L^{2}_{\pi}}^{2}\leq\|\Pi_{r}f_{t}\|_{L^{2}_{\pi}}^{2}+\eta_{r}^{2}.$
Plugging into (4.4), we have
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\mu_{t}\|\pi)\leq-\,\|f_{t}\|_{L^{2}_{\pi}}^{2}+\,\eta_{r}^{2}=-\,\chi^{2}(\mu_{t}\|\pi)+\,\eta_{r}^{2}.$
where we used $\chi^{2}(\mu_{t}\|\pi)=\|f_{t}\|_{L^{2}_{\pi}}^{2}$. Applying the inequality $\chi^{2}(\mu_{t}\|\pi)\geq\mathrm{KL}(\mu_{t}\|\pi)$ to the above inequality, we obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\mu_{t}\|\pi)\leq-\,\mathrm{KL}(\mu_{t}\|\pi)+\ \eta_{r}^{2}.$
Then, by Grönwall’s inequality we have
$\mathrm{KL}(\mu_{t}\|\pi)\leq\mathrm{KL}(\mu_{0}\|\pi)\,e^{-t}+\,\eta_{r}^{2}\,(1-e^{-t}),$
which is exactly (4.2). ∎
Remark 4.4.
Here we assumed sufficient decay or integrability at $\infty$ on the domain $\mathbb{R}^{d}$ so that the “boundary at infinity” term is zero when applying the integration by part.
Remark 4.5.
Proposition 4.3 reveals that the idealized KSWGD (using exact but truncated eigenpairs) inherits the scale-free exponential convergence rate $e^{-t}$ from LAWGD, independent of the Poincaré constant. However, the flow converges to an approximated stationary state with residual error bounded by $\eta_{r}^{2}$. This truncation errors quantifies the steady-state deviation caused by neglecting spectral components beyond the first $r$ modes, and vanishes as $r\to\infty$ by the completeness of the eigenfunction basis, which recovers exact LAWGD.
In practice, the eigenpairs $\{(\lambda_{i},\phi_{i})\}_{i=1}^{r}$ can be estimated from data using EDMD or kernel-EDMD, which gives approximated eigenpairs $\{(\hat{\lambda}_{i},\hat{\phi}_{i})\}_{i=1}^{r}$ and constructs truncated inverse operator $\widehat{\mathcal{K}}_{r}=\sum_{i=1}^{r}\frac{1}{\hat{\lambda}_{i}}\langle\cdot,\hat{\phi}_{i}\rangle_{\pi}\,\hat{\phi}_{i}$ as in (3.2). Therefore, instead of the exact relation $\mathcal{L}\mathcal{K}_{r}=\Pi_{r}$, the data-driven operator satisfies a perturbed relation
$\mathcal{L}\widehat{\mathcal{K}}_{r}f=\Pi_{r}f+\delta_{r}(f),\quad\forall f\in L^{2}_{\pi},$ (4.5)
where $\delta_{r}(f)$ captures the Koopman spectral approximation error.
Assumption 4.6 (Controlled Operator Approximation Error).
There exists a constant $0<\varepsilon_{r}<<1$ such that for all $t\geq 0$,
$|\langle f_{t},\,\delta_{r}(f_{t})\rangle_{\pi}|\;\leq\;\varepsilon_{r}\|f_{t}\|_{L^{2}_{\pi}}^{2}.$
Remark 4.7.
The constant $\varepsilon_{r}$ is directly related to the spectral approximation quality of the Koopman operator.
Theorem 4.8 (Error Bound Analysis).
Let Assumptions 4.1, 4.2 and 4.6 hold. Let $(\mu_{t})_{t\geq 0}$ be the solution to the data-driven KSWGD dynamics with initial distribution $\mu_{0}$:
$\partial_{t}\mu_{t}=\operatorname{div}\big(\mu_{t}\nabla\widehat{\mathcal{K}}_{r}(\mathrm{d}\mu_{t}/\mathrm{d}\pi)\big).$
Then
$\mathrm{KL}(\mu_{t}\|\pi)\;\leq\;e^{-(1-\varepsilon_{r})t}\,\mathrm{KL}(\mu_{0}\|\pi)\;+\;\frac{\eta_{r}^{2}}{1-\varepsilon_{r}}\Big(1-e^{-(1-\varepsilon_{r})t}\Big).$ (4.6)
Proof.
Following the same derivation as Proposition 4.3 up to the dissipation identity in (4.3), but now using the perturbed relation $\mathcal{L}\widehat{\mathcal{K}}_{r}=\Pi_{r}+\delta_{r}$ as in (4.5) instead:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\mu_{t}\|\pi)=-\langle f_{t},\mathcal{L}\widehat{\mathcal{K}}_{r}f_{t}\rangle_{\pi}=-\langle f_{t},\Pi_{r}f_{t}\rangle_{\pi}-\langle f_{t},\delta_{r}(f_{t})\rangle_{\pi}.$ (4.7)
Since $\Pi_{r}$ is self-adjoint and idempotent, $\langle f_{t},\Pi_{r}f_{t}\rangle_{\pi}=\|\Pi_{r}f_{t}\|_{L^{2}_{\pi}}^{2}$. By Assumption 4.6:
$\langle f_{t},\delta_{r}(f_{t})\rangle_{\pi}\leq|\langle f_{t},\delta_{r}(f_{t})\rangle_{\pi}|\leq\varepsilon_{r}\|f_{t}\|_{L^{2}_{\pi}}^{2}.$ (4.8)
Therefore,
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\mu_{t}\|\pi)\leq-\|\Pi_{r}f_{t}\|_{L^{2}_{\pi}}^{2}+\varepsilon_{r}\|f_{t}\|_{L^{2}_{\pi}}^{2}.$ (4.9)
Using the Pythagorean decomposition $\|\Pi_{r}f_{t}\|^{2}=\|f_{t}\|^{2}-\|(I-\Pi_{r})f_{t}\|^{2}$ and Assumption 4.2:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathrm{KL}(\mu_{t}\|\pi)\leq-(1-\varepsilon_{r})\|f_{t}\|_{L^{2}_{\pi}}^{2}+\eta_{r}^{2}=-(1-\varepsilon_{r})\chi^{2}(\mu_{t}\|\pi)+\eta_{r}^{2}.$ (4.10)
Then, applying $\chi^{2}(\mu_{t}\|\pi)\geq\mathrm{KL}(\mu_{t}\|\pi)$ and Grönwall’s inequality gives (4.6). ∎
Remark 4.9.
Theorem 4.8 quantifies the impact of data-driven spectral approximation. Compared to Proposition 4.3, the approximation error $\varepsilon_{r}$ degrades performance in two ways: (i) it slows the exponential convergence rate from $e^{-t}$ to $e^{-(1-\varepsilon_{r})t}$, and (ii) it amplifies the biased error from $\eta_{r}^{2}$ to $\eta_{r}^{2}/(1-\varepsilon_{r})$.
4.2 Discrete-Time Analysis via Approximate Gradient Flow
While our analysis in Proposition 4.3 and Theorem 4.8 establishes exponential convergence for the continuous-time dynamics, practical implementation relies on the discrete-time update rule (Algorithm 1). To rigorously bridge this gap, we adopt the Approximate Gradient Flow (AGF) framework (See [fujisawa2025on] for more details).
The continuous-time dissipation inequality (4.6) derived in the proof of Theorem 4.8 can be restated in the AGF framework, that is, KSWGD is an $(\alpha,\beta)$-approximate $\chi^{2}$-gradient flow with
$\alpha=1-\varepsilon_{r}\quad\text{and}\quad\beta=\eta_{r}^{2},$
where $\alpha$ represents the dissipation rate of the update direction and $\beta$ represents the bias floor. By adapting the discrete-time analysis of [fujisawa2025on] to our $\chi^{2}$-geometry, we can directly translate our continuous-time guarantees to the discrete setting.
Corollary 4.10 (Discrete-Time Linear Convergence).
Consider the discrete KSWGD update $\mu_{t+1}=(\mathrm{I}-h\widehat{v}_{t})\#\mu_{t}$ with step size $h>0$ where $\widehat{v}_{t}=-\nabla\widehat{\mathcal{K}}_{r}(\mathrm{d}\mu_{t}/\mathrm{d}\pi)$ [pmlr-v162-salim22a]. Under the conditions of Theorem 4.8 and assuming $\mu_{t}$ is sufficiently close to $\pi$ (i.e., $\chi^{2}(\mu_{t}\|\pi)\geq\mathrm{KL}(\mu_{t}\|\pi)$), the discrete iterations satisfy:
$\mathrm{KL}(\mu_{t+1}\|\pi)\leq(1-\alpha h)\mathrm{KL}(\mu_{t}\|\pi)+h\beta+\mathcal{O}(h^{2}).$ (4.11)
Iterating this bound gives linear geometric convergence to the noise floor $\beta/\alpha$:
$\mathrm{KL}(\mu_{T}\|\pi)\leq(1-\alpha h)^{T}\mathrm{KL}(\mu_{0}\|\pi)+\frac{\beta}{\alpha}+\mathcal{O}(h).$
Proof.
For any velocity field $v_{t}$, the one-step change in KL satisfies [ambrosio2008gradient, jordan1998variational]
$\mathrm{KL}(\mu_{t+1}\|\pi)-\mathrm{KL}(\mu_{t}\|\pi)\leq-h\langle\nabla\log(\mathrm{d}\mu_{t}/\mathrm{d}\pi),v_{t}\rangle_{L^{2}_{\mu_{t}}}+O(h^{2}).$
From the continuous-time dissipation inequality (4.10) (or its discrete analogue established in Theorem 4.8), the inner product term is bounded by
$\langle\nabla\log(\mathrm{d}\mu_{t}/\mathrm{d}\pi),v_{t}\rangle_{L^{2}_{\mu_{t}}}\leq-\alpha\chi^{2}(\mu_{t}\|\pi)+\beta.$
Thus,
$\mathrm{KL}(\mu_{t+1}\|\pi)\leq\mathrm{KL}(\mu_{t}\|\pi)-h\alpha\chi^{2}(\mu_{t}\|\pi)+h\beta+O(h^{2}).$
Since $\chi^{2}(\mu_{t}\|\pi)\geq 2\mathrm{KL}(\mu_{t}\|\pi)$ holds locally whenever $\mu_{t}$ is sufficiently close to $\pi$, we obtain the desired linear recursion
$\mathrm{KL}(\mu_{t+1}\|\pi)\leq(1-\alpha h)\mathrm{KL}(\mu_{t}\|\pi)+h\beta+O(h^{2}).$
Iterating this inequality $T$ times and summing the resulting geometric series (with ratio $\rho_{h}\coloneqq 1-\alpha h>0$ for $h<1/\alpha$) gives
$\mathrm{KL}(\mu_{T}\|\pi)\leq\rho_{h}^{T}\mathrm{KL}(\mu_{0}\|\pi)+h\beta\sum_{k=0}^{T-1}\rho_{h}^{k}+O(h)\leq(1-\alpha h)^{T}\mathrm{KL}(\mu_{0}\|\pi)+\frac{\beta}{\alpha}+O(h),$
where the last step uses $\sum_{k=0}^{T-1}\rho_{h}^{k}\leq 1/(\alpha h)$. ∎
Remark 4.11.
The boundness of $\mathcal{O}(h^{2})$ in (4.11) is controllable; refer to [fujisawa2025on, pmlr-v162-salim22a] for more details.
Corollary 4.10 bridges the gap between our continuous-time theory and the practical Algorithm (1). Specifically, by identifying the physical time $t=Th$, the discrete geometric decay explicitly recovers the exponential rate derived in Section 4.1, i.e., $(1-\alpha h)^{t/h}\approx e^{-\alpha t}$ as $h\to 0$. Furthermore, the discrete analysis reveals an unavoidable bias $\mathcal{O}(h)$ added to the noise floor which provides a theoretical guide for the efficiency-accuracy trade-off: a smaller step size $h$ reduces bias but requires more iterations. This detail is not obvious in the continuous-time analysis. Finally, we also note that while standard SVGD requires a strong eigenvalue lower bound [fujisawa2025on, Assumption 6] to guarantee a bounded approximation error $\epsilon_{t}$ (a condition often violated by RBF kernels), KSWGD automatically guarantees a bounded error (i.e., constant dissipation rate $\alpha=1-\varepsilon_{r}$) by construction. This is because the rank-$r$ truncation restricts dynamics to the spectral subspace where eigenvalues are strictly lower-bounded, that structurally enables the convergence without external eigenvalue assumptions.
Remark 4.12.
The AGF framework clarifies the theoretical advantage of KSWGD over SVGD. As shown by [fujisawa2025on], standard SVGD suffers from a decaying dissipation rate (i.e., $\alpha_{t}\to 0$), caused by the rapid eigenvalue decay of RBF kernels in RKHS; this inevitably leads to sub-linear convergence. In contrast, KSWGD leverages spectral preconditioning to maintain a constant dissipation rate ($\alpha=1-\varepsilon_{r}>0$) throughout the optimization. This structural difference of preventing the vanishing gradient rate is the fundamental reason why KSWGD achieves linear convergence while SVGD does not in discrete iteration.
5 Feynman-Kac Interpretation of KSWGD
The Koopman operator employed in KSWGD admits a natural interpretation within the Feynman-Kac framework [bertini1995stochastic, karatzas2014brownian, oksendal2003stochastic]. Recall that the general Feynman-Kac formula provides the solution to the following equation:
$\partial_{t}v=\mathcal{A}v-U(x)v,\quad v(x,0)=f(x),$
as the path integral:
$v(x,t)=\mathbb{E}\left[f(X_{t})\exp\left(-\int_{0}^{t}U(X_{s})ds\right)\mid X_{0}=x\right],$
where $U(x)$ acts as a killing (or potential) term and $\mathcal{A}$ is Kolmogorov backward operator (i.e., the infinitesimal generator of the Koopman semigroup associated with the underlying stochastic dynamics as discussed in Section 2.2). The Koopman semigroup $\{\mathcal{T}^{t}\}_{t\geq 0}$ used in KSWGD corresponds to the zero-potential case $U\equiv 0$, namely,
$v(x,t)=(\mathcal{T}^{t}f)(x)=\mathbb{E}[f(X_{t})\mid X_{0}=x].$
What does $U=0$ mean in KSWGD? The potential $U$ in the Feynman-Kac formula serves as a path-dependent weighting (or "killing rate", see AppendixA.4 for more details). It is distinct from the drift potential $V$ that defines the target distribution $\pi\propto e^{-V}$. The drift potential $V$ is already encoded in the Langevin generator $\mathcal{L}=-\mathcal{A}=\nabla V\cdot\nabla-\Delta$. Setting $U=0$ means that we compute unconditional expectations over paths, which is precisely what is needed for sampling the marginal distribution at $t=0$, more specifically, KSWGD targets the distribution of initial conditions $\pi_{0}$ without conditioning on future behavior and the pushforward $(\mathcal{T}^{t})_{\#}\pi_{0}$ gives the marginal distribution at time $t$ without path conditioning. For the Allen-Cahn experiment as shown in Section 6.4, where we sample the unconditional distribution of phase-field configurations and propagate them forward in time, we would have $U=0$.
In contrast, $U\neq 0$ would be appropriate for conditional sampling problems, e.g., sampling initial conditions that lead to a specific terminal state (rare event sampling [weinan2021applied]) or importance-weighted path integrals for computing conditional expectations, in other words, if sampling initial conditions evolve toward a specific target set $B\subset\mathbb{R}^{d}$, then the framework naturally extends to $U\neq 0$ via importance sampling as follows:
$\tilde{\pi}_{0}(x)\propto\pi_{0}(x)\cdot\mathbb{E}\left[\mathbf{1}_{B}(X_{T})\exp\left(-\int_{0}^{T}U(X_{s})ds\right)\mid X_{0}=x\right].$
However, detailed analysis and development of the framework for $U\neq 0$ (e.g., for rare event simulation) are beyond the scope of the current study and represent a promising direction for future research.
Comprehensively, the Feynman-Kac theory extends our proposed method KSWGD to a broader theoretical perspective, which clarifies that the current implementation ($U=0$) is the natural and mathematically appropriate choice for unconditional sampling and prediction tasks (See examples in Section 6.3 and Section 6.4 for more details).
Remark 5.1.
The linear Feynman-Kac formula may appear to conflict with the nonlinear partial differential equation, e.g., Allen-Cahn equation in Section 6.4. However, these two equations describe different objects. The Allen-Cahn equation governs the evolution of the state $u(x,t)$ and is nonlinear in $u$. In contrast, the Feynman-Kac PDE governs the evolution of expected observables $v(x,t)=(\mathcal{T}^{t}f)(x)$, which is linear in $f$. This is precisely the essence of Koopman operator theory: a nonlinear dynamical system acting on states induces a linear operator acting on observables. See [oksendal2003stochastic, Section 8.2] for more details.
6 Experiment
6.1 1-Sphere $S^{1}$
Here we evaluated the proposed KSWGD framework on a 1-dimensional spherical manifold $S^{1}$, targeting a uniform distribution over the unit circle. The data-driven spectral learning utilized $N=500$ training samples $\{x_{i}\}_{i=1}^{N}$ drawn from this target. To construct the necessary time-evolution snapshot pairs $(x_{i},\hat{y}_{i})$ without knowing the potential $V$, we simulated dynamics over a short interval $\Delta t=0.05$, approximating the drift term $\nabla\log\pi$ via Kernel Density Estimation (KDE) [parzen1962estimation]. To strictly enforce the geometric structure, Langevin SDE updates were projected onto the local tangent space and renormalized to unit length at each step. Based on these pairs, the Koopman operator was approximated via Kernel-EDMD [kevrekidis2016kernel] using both a Polynomial kernel of degree 10 (see Figure 1(a)) and a Gaussian RBF kernel (see Figure 1(b)), using Tikhonov regularization with $\gamma=10^{-6}$ for numerical stability.
For the generative process, we initialized $M=700$ particles concentrated at the top of the 1-sphere; specifically, $y>0.7$ depicted in red dots, and evolved them according to the learned Koopman spectral gradients (i.e., KSWGD) for $T=1000$ iterations with a step size of $h=2$. As shown in Figure 1, KSWGD successfully push the particles out and uniformly cover the manifold, as depicted in the purple circles. In contrast, a comparative analysis using the Diffusion Map Particle System (DMPS) as a baseline [li2025diffusion] under identical settings fails to effectively cover the manifold. The result is presented in Figure 5 of Appendix A.1.1.
(a) Kernel: RBF
(b) Kernel: Polynomial
Figure 1: 1-Sphere $S^{1}$ example using KSWGD with Kernel-EDMD.
6.2 Quadruple Well System
Here we consider a metastable two-dimensional quadruple-well potential $V(x,y)=(x^{2}-1)^{2}+(y^{2}-1)^{2}$ characterized by four local minima at $(\pm 1,\pm 1)$. The objective is to sample from the Boltzmann distribution $\pi(x)\propto\exp(-V(x))$ of the associated overdamped Langevin dynamics without explicit knowledge of $V$. To construct the training dataset, we generate a trajectory from the stationary distribution via MCMC sampling and evolve each state forward using an Euler–Maruyama discretization with a time step $\Delta t=0.1$. This produces 2,500 consecutive time-series pairs $(X_{t},X_{t+\Delta t})$. The Koopman operator is then approximated using SDMD [10.1063/5.0283640] with a neural network-based dictionary.
For the generative process via KSWGD, we initialize 500 particles from the Langevin dynamics and discard those already located inside any of the four wells using an exclusion radius of 0.4, retaining particles primarily in the transition regions. We then apply KSWGD for $T=1000$ iterations with a step size of $h=1$. As illustrated in Figure 2, the red dots denote the initial particle positions, while the purple circles indicate their final positions after 1000 steps. It can be observed that each particle is successfully transported to and stops at the "bottom" of the well nearest to its initialization. Notably, the final positions do not lie exactly at the theoretical minima, which is consistent with the bias term $\eta_{r}^{2}/(1-\varepsilon_{r})$ characterized in Theorem 4.8. A detailed comparison with DMPS [li2025diffusion] is provided in Figure 6 and Figure 7 of Appendix A.1.2.
(a) all partticles
(b) 15 partticles
Figure 2: Quadruple potential well example using KSWGD with SDMD.
6.3 MNIST
In this experiment, we evaluate the effectiveness of KSWGD for high-dimensional image generation using the MNIST handwritten digit dataset, which consists of 60,000 training samples and 10,000 test samples of 28×28 grayscale images. A CNN-based autoencoder is trained to map the original 784-dimensional image space to a 6-dimensional latent space. In this latent space, we construct a Koopman operator approximation using EDMD with a Mini-Batch Dictionary Learning scheme [mairal2009online], where the sparsity regularization coefficient is set to $\alpha=10^{-3}$. Dictionary features are learned directly from the latent representations $z_{i}$ corresponding to training samples $x_{i}$, and the spectral decomposition of the Koopman operator is obtained via a generalized eigenvalue problem. We set the regularization parameter to $10^{-3}$ and apply an eigenvalue truncation threshold of $10^{-6}$ when forming the spectral expansion used by KSWGD.
During the sample generation stage, KSWGD operates entirely in the latent space using 64 particles initialized from a standard Gaussian distribution $\mathcal{N}(0,I)$. Particles are iteratively updated with a fixed step size of 0.1 over 500 iterations, progressively evolving toward the target distribution under KSWGD dynamics. After decoding the final particle states through the decoder network, the generated samples exhibit clearly recognizable digit structures, as shown in Figure 3, which demonstrates that KSWGD can model complex high-dimensional data distributions. Additional experiments with multiple random seeds $\{1,2,3\}$ and step sizes $\{0.05,0.2\}$ are presented in Appendix A.1.3, which further confirm the numerical stability and robustness of the method as shown in Figure 8. In contrast, the DMPS baseline fails to produce meaningful digit-like patterns, as illustrated in Figure 9.
Remark 6.1.
In the MNIST experiment, we employ a standard autoencoder to obtain a latent representation for dimension reduction purpose; however, the quality of this latent space naturally depends on different choices of the architecture and hyperparameter settings, which we did not extensively tune for. Investigating optimal latent space construction or alternative dimension reduction techniques is beyond our scope. Our focus is on validating KSWGD’s sampling capability given a reasonable latent representation.
Figure 3: MNIST example using KSWGD with EDMD (Dictionary Learning).
6.4 Stochastic Allen-Cahn Equation
Here we evaluate the predictive capability of KSWGD on the stochastic Allen-Cahn equation [allen1972ground, steinbach2009phase], which models phase separation kinetics driven by thermal fluctuations and serves as a benchmark for testing generative methods on SPDEs. The system is governed by:
$du=\left[D\nabla^{2}u-\epsilon^{-2}u(u^{2}-1)\right]dt+\sigma dW_{t},$
where $u(x,t)$ is the phase field, $D$ is the diffusion coefficient, and $W_{t}$ represents the standard Wiener process multiplied by noise intensity $\sigma$. The parameter $\epsilon$ defines the width of the interfacial transition layers. A critical property of this system is its stiffness, more specifically, as $\epsilon\to 0$, the reaction term $\epsilon^{-2}u(u^{2}-1)$ becomes large, which forces rapid transitions towards the stable equilibria $u=\pm 1$. This necessitates a very fine temporal resolution of $\Delta t$ to maintain numerical stability and accurately resolve the fast interface dynamics, making data-driven modeling particularly challenging due to the high-dimensional spatiotemporal correlations.
In this experiment, we validate KSWGD’s capability for learning these spatiotemporal distributions. We simulate the equation on a $128\times 128=16,384$ grid with time step $\Delta t=10^{-5}$, $D=0.001$, $\epsilon=0.01$, and $\sigma=1.4142\approx\sqrt{2}$. The small interface width parameter $\epsilon$ drives rapid reaction dynamics, which necessitates fine temporal resolution to accurately capture the phase separation process. We generate 300 independent realizations with snapshots recorded at $t\in\{0,0.0001,0.0002,0.0005\}$. Each 16,384-dimensional snapshot is compressed to an 16-dimensional latent space via a fully-connected autoencoder. Using EDMD with 2nd-order polynomial features, we construct the Koopman matrix $K$ in latent space from paired data at $t_{0}=0$ and $t_{1}=0.0001$, with regularization parameter $10^{-4}$. During the KSWGD sampling phase, 150 particles are initialized in the 8-dimensional latent space from a Gaussian distribution which is rescaled to match the statistics of $Z_{t_{0}}\coloneqq[z^{(1)}_{t_{0}},z^{(2)}_{t_{0}},...,z^{(150)}_{t_{0}}]^{T}\in\mathbb{R}^{150\times 8}$ (i.e., the latent encodings of the 150 initial snapshots at $t_{0}=0$), and evolved over 800 iterations with step size $h=0.03$ to generate samples matching the latent distribution at $t_{0}=0$. Subsequently, we apply the Koopman matrix $K$ successively in latent space to predict future distributions: one application $K^{1}$ predicts $t=0.0001$, three applications $K^{2}$ predict $t=0.0002$, and five applications $K^{5}$ predict $t=0.0005$. Notice that the snapshots at $t_{2}=0.0002$ and $t_{5}=0.0005$ are not used in training, and serve purely to validate the Koopman operator’s predictive capability beyond the training time (See Appendix A.3 for more details). Figure 4 shows that the decoded KSWGD samples are highly consistent with the ground truth in both spatial and statistical features, which reflects the operator’s ability to capture the essential distributional dynamics underlying the physical trajectory evolution.
Moreover, we also present comparison tests with other baseline methods including Diffusion Modeling (DDPM) [ho2020denoising, nichol2021improved], VAE [kingma2013auto], Normalizing Flows [dinh2017realnvp] and GAN [gulrajani2017improved] in Figure 12, as well as other tests with different parameter $\epsilon=0.01$ in Figure 11 and Figure 12 of Appendix A.1.4. Notice that smaller $\epsilon$ creates sharper phase interfaces (i.e., approaching the sharp-interface limit) with faster dynamics, which requires finer spatial resolution to capture high-frequency features and smaller timesteps for numerical stability. The details of the settings of these methods are discussed in Appendix A.5.
Figure 4: Allen-Cahn equation example using KSWGD with EDMD (Polynomial): $\epsilon=0.01$
7 Conclusion
In this work, we have introduced a training-free generative modeling framework called KSWGD that bridges Koopman operator theory with Wasserstein gradient flows. By the fundamental connection between the Koopman generator and the Langevin generator, our method enables data-driven spectral construction of the kernel as in LAWGD without access to the target potential. The theoretical analysis establishes that KSWGD achieves scale-free exponential convergence with quantifiable error bounds depending on the rank of spectral truncation and data-driven approximation quality. A key structural advantage over SVGD is the maintenance of a constant dissipation rate through spectral preconditioning, which prevents the vanishing gradient phenomenon that leads to sublinear convergence in discrete setting. The Feynman-Kac interpretation further extend KSWGD to a broader theoretical framework, with the zero-killing rate case ($U\equiv 0$) being the natural choice for unconditional sampling tasks. Experimental results on problems ranging from low-dimensional compact manifolds and metastable systems to high-dimensional image generation and SPDE validate the method’s practical effectiveness and robustness. Extending the framework to conditional sampling via non-zero killing rates and larger-scale applications remains an interesting direction for future work.
Acknowledgments
Authors are grateful to Ben Gao, Hanru Bai, Zhicheng Jiang, Qiao Sun, Sinho Chewi, Vladimir Kostic, Yunan Yang, Micah Warren, Maria Oprea, Zhongwei Shen, Yao Li, for all kinds of valuable discussion and comments. Y.X. and I.I. acknowledge support from JST CREST Grant Number JPMJCR24Q1 including AIP challenge program, Japan.
Appendix A Experiments
A.1 Comparison to Other Methods
Here we are going to show comparison tests between KSWGD and other baseline methods DMPS [li2025diffusion], VAE [kingma2013auto], Diffusion Modelling [nichol2021improved].
A.1.1 1-Sphere $S^{1}$
We also compare KSWGD and DMPS on the 1-dimensional spherical manifold $S^{1}$, as shown in Figure 5. Under the same experimental settings, DMPS fails to reach full convergence to the target distribution, even when the number of iterations or the step size is moderately increased. This further highlights the advantage of KSWGD’s spectral preconditioning in sustaining an effective gradient flow throughout the optimization.
Figure 5: 1-Sphere $S^{1}$ example using DMPS with different step size $h$.
A.1.2 Quadruple Potential Well System
Here, we further apply DMPS to the quadruple-well example. Compared with KSWGD, DMPS converges significantly more slowly: many particles remain outside the wells at iteration 1000, and the method becomes stable at 3000 iterations, as shown in Figure 6. To quantify this difference, we measure two metrics for both algorithms: (i) the percentage of particles that have entered one of the wells, and (ii) the average $\text{Movement Rate}=\frac{1}{N}\sum_{i=1}^{N}\lVert x_{i}^{(t+1)}-x_{i}^{(t)}\rVert$, which reflects the averaged particle velocity and serves as a direct indicator of dynamical stability. A movement rate near zero indicates that particles have effectively stopped moving. Under the convergence criterion “movement rate $\leq 0.01$ and well coverage $\geq 95\%$,” KSWGD reaches practical convergence in fewer than 500 iterations, whereas DMPS requires roughly 2000 iterations to stabilize, as shown in Figure 7. This contrast highlights the significantly faster convergence of KSWGD relative to DMPS.
(a) all partticles
(b) all partticles
Figure 6: Quadruple potential well example using DMPS at 1000th and 3000th iteration step.
Figure 7: Comparison: DMPS vs KSWGD-SDMD.
A.1.3 MNIST
In this section, we provide additional results to assess the algorithm’s sensitivity to hyperparameter variations. Specifically, we illustrate the generation performance under different random seeds $\{1,2,3\}$ and step sizes $\{0.05,0.2\}$ in Figure 8. We also include the DMPS outputs for comparison in Figure 9.
Figure 8: Extra MNIST examples using KSWGD with EDMD.
Figure 9: MNIST example using DMPS.
A.1.4 Stochastic Allen-Cahn Equation
Here we also show the comparison tests with other baseline methods Diffusion Modelling (DDPM), VAE, Normalizing Flows and GAN in Figure 11 and Figure 12. Moreover, we show $\epsilon=0.02$ case, as shown in Figure 4 and Figure 10.
Figure 10: Comparison: KSWGD vs DM, VAE, Normalizing Flows and GAN, $\epsilon=0.01$.
Figure 11: Allen-Cahn equation example using KSWGD with EDMD (polynomial dictionary): $\epsilon=0.02$
Figure 12: Comparison: KSWGD vs DM, VAE, Normalizing Flows and GAN, $\epsilon=0.02$.
A.2 KDE-Based Score Estimation
The score function $\nabla\log\pi(x)$ is estimated using a Gaussian kernel density estimator as following:
First we compute pairwise squared distances:
$D_{ij}^{2}=\|x_{i}-x_{j}\|^{2},\quad i,j=1,\ldots,n.$
Then we select bandwidth by median heuristic: $h=\sqrt{\text{median}(D^{2})}$. Next, we compute the Gaussian kernel weights:
$W_{ij}=\exp\left(-\frac{\|x_{i}-x_{j}\|^{2}}{2h^{2}}\right),$
and normalize weights $w_{ij}=\frac{W_{ij}}{\sum_{k=1}^{n}W_{ik}}$. Then we compute weighted mean $\bar{x}_{i}=\sum_{j=1}^{n}w_{ij}x_{j}$. The score function estimation is give as follows:
$\nabla\log\hat{\pi}(x_{i})=\frac{\bar{x}_{i}-x_{i}}{h^{2}}=\frac{1}{h^{2}}\left(\frac{\sum_{j=1}^{n}W_{ij}x_{j}}{\sum_{j=1}^{n}W_{ij}}-x_{i}\right).$
This estimator approximates the gradient of the log-density by computing the direction from each point $x_{i}$ toward its local kernel-weighted centroid, scaled by the squared bandwidth $h^{2}$.
A.3 Koopman Operator Prediction in Latent Space
To predict the time evolution of latent dynamics in the Stochastic Allen-Cahn equation, we employ an Extended Dynamic Mode Decomposition (EDMD) approach with polynomial dictionary functions. Let $z\in\mathbb{R}^{d}$ denote the latent vector. The central idea is to lift the latent representation into a higher-dimensional feature space where the dynamics become approximately linear.
We define a polynomial feature map $\Phi:\mathbb{R}^{d}\to\mathbb{R}^{N_{K}}$ that includes all monomials up to degree $p$:
$\Phi(z)\coloneqq[\underbrace{1}_{\text{constant}},\underbrace{z_{1},\ldots,z_{d}}_{\text{coordinates}},\underbrace{z_{1}^{2},z_{1}z_{2},\ldots,z_{1}^{p},\ldots,z_{d}^{p}}_{\text{higher-order terms up to degree $N_{K}$}}]$
Crucially, the original latent coordinates $z_{1},\ldots,z_{d}]$ appear explicitly as second through $(d+1)$-th entries in this dictionary $\Phi(z)$. Given paired trajectory data $\{(z_{i}^{(0)},z_{i}^{(\Delta t)})\}_{i=1}^{N}$ sampled at times $t=0$ and $t=\Delta t$, we estimate the finite-dimensional Koopman operator $K\in\mathbb{R}^{N_{K}\times N_{K}}$ by solving the least-squares problem:
$\Phi(z^{(\Delta t)})\approx\Phi(z^{(0)})\,K.$
For multi-step prediction, we first compute its lifted representation $\Phi(z_{t})$ starting from a latent state $z_{t}$, then propagate forward in the feature space via $\Phi_{t+\Delta t}=\Phi(z_{t})K$. To recover the predicted latent state, we simply extract the $z_{1},\ldots,z_{d}]$ part from $\Phi_{t+\Delta t}$, which gives the updated coordinates directly. This operation is repeated to generate trajectories over arbitrary time steps, which enables the Koopman operator to capture the nonlinear latent dynamics through the polynomial dictionary.
A.4 Probabilistic Interpretation of the Killing Rate
In the linear Feynman-Kac formula, we have
$u(x,t)=\mathbb{E}\left[f(X_{t})\exp\left(-\int_{0}^{t}U(X_{s})\,ds\right)\mid X_{0}=x\right],$
where the function $U(x)$ is related to the killing rate. This terminology arises from a probabilistic interpretation, more specifically, consider a particle following the stochastic trajectory $X_{s}$. At each position $x$, the particle is "killed" (removed from the system) at rate $U(x)$. The exponential term $\exp\left(-\int_{0}^{t}U(X_{s})\,ds\right)$ then represents the survival probability of the particle along its path up to time $t$. When $U(x)>0$, paths passing through regions of high $U$ are down-weighted; when $U\equiv 0$, all paths contribute equally.
In our KSWGD framework, we set $U\equiv 0$. This is not an approximation but rather the mathematically appropriate choice for our task. With $U=0$, the Feynman-Kac formula reduces to the standard Koopman semigroup $\mathcal{K}^{t}f(x)=\mathbb{E}[f(X_{t})\mid X_{0}=x]$, which computes unconditional expectations over all paths. This is precisely what is required for sampling the marginal distribution at the initial time, and predicting future marginal distributions via Koopman propagation.
A non-zero killing rate $U\neq 0$ would correspond to a different class of problems, such as conditional sampling or rare event simulation, where one seeks initial conditions whose trajectories satisfy certain path-dependent constraints. For instance, if $U$ penalizes high interface energy, the resulting distribution would favor initial conditions that lead to low-energy configurations. Such extensions, while natural within the Feynman-Kac framework, lie outside the scope of the present work.
A.5 Baseline Comparison Details for Stochastic Allen-Cahn Equation
To provide a fair and computationally tractable comparison, we first train a shared fully-connected autoencoder that maps the high-dimensional $128\times 128=16384$-dimensional spatial fields to an 8-dimensional latent representation through an encoder network ($16384\to 512\to 128\to 8$) and reconstructs them via a symmetric decoder. All generative methods, including the baselines, then operate in this shared 8-dimensional latent space rather than the original image space. This design choice ensures that all methods benefit from the same dimensionality reduction and that differences in performance reflect the generative modeling capabilities rather than representation quality. For each experiment, we generate 300 independent SPDE realizations with snapshots recorded at $t\in\{0,0.0001,0.0002,0.0005\}$, using simulation parameters $\Delta t=10^{-5}$, $D=0.001$, and $\sigma=\sqrt{2}$. All baseline methods are trained for 150 epochs using the Adam optimizer, and each generates 100 samples per time point for evaluation. It is worth noting that KSWGD learns a Koopman operator from $t_{0}\to t_{1}$ pairs and predicts future distributions via powers of $K$, whereas each baseline model is trained separately on the ground-truth samples available at each time point.
Diffusion Model (DDPM).
We implement a latent-space denoising diffusion probabilistic model. The model consists of an MLP-based noise predictor with hidden dimension 256 and a sinusoidal time embedding of dimension 64. During training, we use a linear beta schedule from $10^{-4}$ to $0.02$ over 200 diffusion steps, and the network learns to predict the noise added to latent codes at each diffusion timestep. For sampling, we initialize from pure Gaussian noise $z_{T}\sim\mathcal{N}(0,I)$ in the 8-dimensional latent space and iteratively denoise for 200 steps using the learned predictor, then decode the resulting latent code back to physical space via the shared autoencoder.
Variational Autoencoder (VAE).
The VAE operates as a second-level encoder-decoder within the shared latent space, mapping the 8-dimensional AE latent codes to an even smaller 4-dimensional VAE latent space. The encoder and decoder each consist of two hidden layers with dimension 128 and ReLU activations. We optimize the evidence lower bound (ELBO) with a KL divergence weight $\beta=0.1$ to balance reconstruction quality and latent space regularity. During generation, we sample from the 4-dimensional standard Gaussian prior $w\sim\mathcal{N}(0,I)$ and decode to obtain new 8-dimensional latent codes, which are then mapped back to physical space via the shared AE decoder.
Normalizing Flows (RealNVP).
We implement a latent-space RealNVP-style normalizing flows with 4 affine coupling layers, each parameterized by an MLP with hidden dimension 128. The coupling layers use alternating binary masks to partition the 8 latent dimensions, and the scaling outputs are stabilized via a tanh nonlinearity. The model is trained by maximizing the exact log-likelihood of the latent codes using the change-of-variables formula. For sampling, we draw from a standard Gaussian base distribution $w\sim\mathcal{N}(0,I)$ and apply the learned inverse transformation $z=f_{\theta}^{-1}(w)$ to obtain samples in the AE latent space, followed by decoding to physical space.
GAN (WGAN-GP).
We employ a latent-space Wasserstein GAN with gradient penalty for stable adversarial training. The generator maps 16-dimensional Gaussian noise to 8-dimensional latent codes through an MLP with hidden dimension 256, while the critic uses a similar architecture to distinguish real from generated latent codes. We use 5 critic updates per generator update with a gradient penalty coefficient $\lambda=10.0$, and train with Adam using learning rate $10^{-4}$ and momentum parameters $(\beta_{1},\beta_{2})=(0.5,0.9)$. Samples are generated by drawing noise $\xi\sim\mathcal{N}(0,I)$ and passing through the trained generator to obtain latent codes, which are then decoded to physical space.